OpenStudy (anonymous):

I need to find all real solutions of (X^2 + 7)^3/2 = 64 I can get to [(x^2 + 7)^3/2]^2/3 = 64^2/3 then I'm stuck please help

OpenStudy (anonymous):

64=4^3 64^(2/3)=(4^3)^(2/3) =4^2=16 so x^2+7=16

OpenStudy (anonymous):

proceed now lolll its easy now

OpenStudy (anonymous):

Got it! Thanks